• The value of m for which 2x+3y=11 , y=mx+3 , 2x-4y=-24 has common solution is?
1
Expert's answer
2013-02-13T09:32:28-0500
First, we find the solution for (1) 2x+3y=11
(2) 2x-4y=-24
(1) - (2):
2x-2x+3y-(-4y) = 35 => 7y=35
y=5
From the first:
2x+3*5=11 => 2x=-4
x=-2
Therefore y=mx+3 also has solution (-2, 5)
5=m*(-2) +3
m=-1
Answer: m=-1.